OpenStudy (anonymous):

log(x+5)-log(x-3)=log5 want to get checked! logx-3/x+5=log5 drop the log . x-3/x+5=5 (x+5)*(x-3/x+5)=5(x+5) x-3=5x+25 -28=4x x=-7

OpenStudy (amistre64):

consider - to be under

OpenStudy (amistre64):

- log(n) = /n in the setup

OpenStudy (anonymous):

logx-3/x+5=log5 Is this step correct?

Parth (parthkohli):

Just check it by inputting the following into the calculator. \( \color{Black}{\Rightarrow (\log (-7) + 5) - (\log(-7) - 3)}\) and \(\log5\) Check if they give the same result.

OpenStudy (anonymous):

I think it should have been logx+5/x-3=log5

OpenStudy (anonymous):

make sure it's non negative too.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \log(-2) - \log(-10) = ?}\)

Parth (parthkohli):

Doesn't look like we have a definite number ;)

OpenStudy (anonymous):